# Transforming Strategic Games into Biform Games: Applications in Allocation Mechanisms and Green Technology Investment

### v1

@@ Abstract
As Aumann stated, cooperation and non-cooperation are different ways of viewing the same game, with the main difference being whether players can reach a binding cooperative agreement. In the real world, many games often coexist competition and cooperation. Based on the above reasons, we propose a method to transform strategic games into a biform game model, which retains the characteristics of cooperative games while considering the ultimate goal of players to maximize their own interests. Furthermore, based on this biform game model, we analyze the impact of two different distribution methods, namely marginalism and egalitarianism, on the game results. As an application, we analyze how food producers seek maximum profits through cooperative pricing.
@@ 1 Introduction
This paper mainly includes two parts: the first part is to establish a research method of transforming from a strategic game (players can choose a cooperative game or a non-cooperative game) to a biform game; in the second part, we illustrate the method and sentencing game model proposed in the first part through two examples, and utilize this conversion method and biform model to analyze how food producers seek maximum profits through cooperative pricing..
@@
Let’s turn our attention to cooperative game theory. The most important thing in cooperative game theory is the solution to the cooperative game, which is to answer the question: When we assume that the result of cooperation is the benefit of the grand coalition, how should the result of cooperation be distributed among the players in a reasonable way? The rationality mentioned here includes fairness, the enthusiasm to encourage players to form alliances, and a certain degree of stability (the coalition is not easily destroyed). From a large body of literature, it is not difficult to find that the solutions of cooperative games are inseparable from the two logics of marginalism and egalitarianism. The Shapley value is a typical example of marginalism, while the equal division rule is a typical example of egalitarianism [14, 15, 16, 17, 18]. The Shapley value does not allocate anything to non-productive players (those who do not make a contribution), while the equal allocation rule does not consider the contribution rate of players to the coalition’s revenue. However, neither of these two methods is completely consistent with the form of resource or outcome distribution in practical problems. When distributing the fruits of cooperation, we not only give incentives to players with high contributions, but also pay attention to the common sharing of the fruits of cooperation. What impact will marginalism and egalitarianism have on the final result in the distribution? How should our distribution method seek a balance between the two for specific problems? Or, what kind of distribution plan should we formulate in order to achieve the goal? These are the questions that the first part of this chapter attempts to answer. This part will systematically study the impact of the distribution of coalition benefits based on the biform game derived from the strategic game, and explore a method to analyze how the distribution of coalition utility affects strategic behavior.
@@ 2.1 Problem Statement
Suppose the strategic game is $\left({N,X,f}\right)$. Here $N=\left\{{1,\cdots,n}\right\}$ is the set of players, $X_{i}$} represents the strategy set of participant $i$, $X=\mathop{\prod}\nolimits_{i\in N}{X_{i}}$ is the strategy combination set, ${x_{i}}\in{X_{i}}$ represents the strategy selected by participant $i$, $\left({{x_{i}},{x_{-i}}}\right)\in{\text{X}}$ represents the strategy combination formed after each participant $i$ selects $x_{i}$}, and $x_{-i}$} represents the strategy selection component of all players except $i$. $f=\left({{f_{1}},\cdots,{f_{n}}}\right)$ is a vector of payoff functions, where ${f_{i}}:X\to R$ is the payoff function of player $i$. In particular, when $N=\left\{{1,2}\right\}$ and ${X_{1}}={S_{1}}$ and ${X_{2}}={S_{2}}$ are finite sets, then the two-person finite game $\left({\left\{{1,2}\right\},\left({{S_{i}}}\right),\left({{f_{i}}}\right)}\right)$ can be represented by a table, that is, the two payoff matrices are combined into one table.
When a strategic game is considered a non-cooperative game, we still use its notation, but assign the strategy set payoff function to each participant, namely, $\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$. In this scenario, each participant seeks to maximize their own payoff. A state of equilibrium, known as the Nash equilibrium the central solution concept in non-cooperative game theory is reached when no player can obtain a higher payoff by unilaterally changing their strategy.
@@
In non-zero-sum situations, we have ${f_{1}}\left({x_{1}^{*},x_{2}^{*}}\right)>v\left({\left\{1\right\}}\right)$ and ${f_{2}}\left({x_{1}^{*},x_{2}^{*}}\right)>v\left({\left\{2\right\}}\right)$. For example, the Battle of the Sexes game, the Hawk-Dove game, and the Wise Pigs game (also known as the free-rider game) all satisfy this property. Note that ${f_{1}}\left({x_{1}^{*},x_{2}^{*}}\right)$ and ${f_{2}}\left({x_{1}^{*},x_{2}^{*}}\right)$ are the payoffs under non-cooperation, while $v\left({\left\{1\right\}}\right)$and $v\left({\left\{2\right\}}\right)$are the payoffs under cooperation. Considering the opponent’s confrontation actually reduces the payoffs of the players, which seems logically unreasonable. Therefore, it is unreasonable to assume that players other than $S$ choose a confrontational strategy.
@@
Furthermore, transforming strategic games into cooperative games focuses on cooperation between players, but fails to fully reflect their pursuit of self-interest. To overcome these limitations, we propose a biform game model derived from strategic games. This model assigns a coalition value to each strategy combination, regardless of the strategies chosen by players within or outside the coalition. Cooperation within the coalition primarily manifests itself in utility transfer and agreement on strategy selection.
@@ 2.2 Model Building
Similar to expressions (1), (2) and (3), ${u_{f}}\left({S,x}\right)$ represents the payoff of coalition $S$ at the strategy combination $x$, referred to as the value of coalition $S$ at $x$, which is the total payoff of the players in coalition $S$.
As mentioned earlier, the characteristic function defined by (1), (2) and (3) must be either the maximum or minimum value on $X_{S}$} or the corresponding $X_{-S}$}. By defining the coalition value according to (4), there is no need to worry about whether to maximize or minimize it, as with $v$, $v_{H}$}, and $v_{M}$}. A way to circumvent this problem is to give, for each strategy combination $x\in X$, the value ${u_{f}}\left({S,x}\right)$ of the coalition $S$ at $x$, then, ${u_{f}}\left({\cdot,x}\right):{2^{X}}\to R$ is the characteristic function of the strategy combination $x$, or the coalition function associated with the strategy combination $x$. For each $x\in X$, ($N,{u_{f}}\left({\cdot,x}\right)$) constitutes a cooperative game.
According to the biform game model proposed by Brandenburger and Stuart [4], cooperative games with strategy combinations $x$ constitute biform games. Biform games first appeared in concrete examples ([2] and Hart and Moore [3]), and the concept of biform games and their solutions was first introduced by Brandenburger and Stuart [4].
@@
Here $\delta\left({S,x}\right)\geqslant 0$ indicates that in addition to the cumulative benefits of the players, in some cases, coalitions can bring additional benefits, such as collaborative effects, government subsidies, and cost savings after supply chain members form a coalition, as well as logistics and warehousing costs. Therefore,
@@
We still call ${v_{f}}:{2^{N}}\times X\mapsto R$ the characteristic function with the strategy combination $x$, or the coalition function with the strategy combination $x$. In this case, $u_{f}$} becomes a special case of $v_{f}$}, that is, the case where the additional benefit $\delta\left({\cdot,x}\right)=0$.
Furthermore, considering that when players cooperate, in addition to the utility transfer and collaborative effects between players, another possibility is to restrict the scope of strategy choices through agreement. From the perspective of the strategy combination set $X$, this means adding constraints or limiting strategy choices to a subset $X^{C}$} of $X$. $X^{C}$} is called the collaboration set of the strategic game $\left({N,X,f}\right)$, or simply the collaboration set. Based on the above analysis, the following definition is given.
Definition 2.1 For a given strategic game $\left({N,X,f}\right)$, ${v_{f}}:{2^{N}}\times X\mapsto R$ is the characteristic function with the strategy combination $x$, ${X^{C}}\subset X$ is the cooperation set, then $\left({N,{v_{f}},{X^{C}}}\right)$ constitutes a biform game, which is called a cooperative game and a biform game derived from $\left({N,X,f}\right)$.
Remark 2.1 (1) For every $x\in X$, $\left({N,{v_{f}}\left({\cdot,x}\right)}\right)$ is the cooperative game with the strategy combination $x$, also called the cooperative game in $x$ derived from $\left({N,X,f}\right)$.
@@
(3) ${X^{C}}\subset X$ represents a mutual agreement among players to appropriately limit the range of strategic options, including government intervention in prices. A major role of the Organization of the Petroleum Exporting Countries (OPEC) is to restrict oil prices or production for its members. Recently, the state’s efforts to combat involution in certain industries have also involved interventions in production capacity and prices.
For the biform game $\left({N,{v_{f}},X}\right)$ derived from the strategic game $\left({N,X,f}\right)$, the first stage is a non-cooperative game, primarily describing the strategic choices of the players. This involves the interactive process of each participant selecting different strategies $x_{i}$} (forming a strategy combination $x$) to maximize their payoff. This also reflects the mutual influence of strategy choices as players pursue maximum payoff. The second stage is concerned with each cooperative game $\left({N,{v_{f}}\left({\cdot,x}\right)}\right)$, which reflects the formation of alliances among players and the distribution of cooperative gains among them. It determines the payoff value each participant receives when a strategy profile is realized, and also serves as the payoff function for the first-stage non-cooperative game. Transforming a strategic game into a biform game represents a novel analytical approach. It not only addresses the shortcomings of directly converting a strategic game into a cooperative game such as relying on max-min methods but also constitutes a new attempt to consider competition and cooperation simultaneously.
@@
Let $N=\left\{{1,2}\right\}$ denote the set of players (two herders), ${X_{1}}={X_{2}}=\left\{{C,NC}\right\}$ denote their strategy set, $X={X_{1}}\times{X_{2}}$ denote the strategy combination set, ${f_{1}}\left(x\right),{f_{2}}\left(x\right)$ denote the payoff functions of herders 1 and 2 respectively, and each $f\left(x\right)=\left({{f_{1}}\left(x\right),{f_{2}}\left(x\right)}\right)$ denote a payoff vector, where
@@
Let $\left({N,{v_{f}}\left({\cdot,x}\right)}\right)$ be the cooperative game derived from $\left({N,{v_{f}},X}\right)$. For each player $i$, $S\subset N$ is a coalition and $i\notin S$, then define $i$ as the marginal contribution of the coalition $S$:
@@
$m_{i}^{{v_{f}}}:{2^{N}}\times X\to R$ is called the marginal contribution function with respect to $v_{f}$}.
Definition 2.2 Let $\left({{\text{N}},{v_{f}},X}\right)$ be a biform game derived from $\left({N,X,f}\right)$. The marginal contribution function $m_{i}^{{v_{f}}}$is called payoff-dominant if for every $i\in N$ and $x,y\in X$, ${f_{i}}\left(x\right)>{f_{i}}\left(y\right)$, it follows that $m_{i}^{{u_{f}}}\left({S,x}\right)>m_{i}^{{u_{f}}}\left({S,y}\right)$.
Remark 2.2 If for every $i\in N$ and $x,y\in X$, ${f_{i}}\left(x\right)>{f_{i}}\left(y\right)$, it follows that $m_{i}^{\delta}\left({S,x}\right)>m_{i}^{\delta}\left({S,y}\right)$, then $m_{i}^{{v_{f}}}$ is payment-dominated. This means that the larger the payoff to the participant, the larger the value of the additional benefit distributed. In particular, if $\delta\left({\cdot,x}\right)=c$, it is clear that $m_{i}^{{v_{f}}}$ is payment-dominated, and thus it is clear that $m_{i}^{{u_{f}}}$ is payment-dominated.
2.3 Distribution of Cooperative Outcomes and Biform Game Solutions
Let $\left({N,X,f}\right)$ be the strategic game, and $\left({N,{v_{f}},X}\right)$ be the biform game derived from it. Before presenting the solution to the biform game, we first discuss an important issue: the distribution of coalition payoffs during the cooperative phase. This determines the important properties of the biform game solution. When returning to the first stage of the non-cooperative game, the Nash equilibrium will be primarily influenced by the strategic choices of the players, whose goal is to maximize their respective payoffs, which are determined by the allocation function.
Suppose the vector ${a^{{v_{f}}}}\left(x\right)=\left({a_{1}^{{v_{f}}}\left(x\right),\cdots,a_{n}^{{v_{f}}}\left(x\right)}\right)$ is the solution of the cooperative stage $\left({N,{v_{f}}\left({\cdot,x}\right)}\right)$ of the cooperative game. Each $a_{i}^{{v_{f}}}\left(x\right)$ is the distribution of the grand coalition payoff ${v_{f}}\left({N,x}\right)$ to the payoff of participant $i$. We call the vector-valued function ${a^{{v_{f}}}}:X\mapsto{R^{n}}$ an allocation function of $v_{f}$}. The function $a_{i}^{{v_{f}}}\left(x\right)$ gives a new payoff to player $i$, which replaces the initial payoff value ${f_{i}}\left(x\right)$ in the strategic game. The non-cooperative game $\left({N,\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$ constitutes the non-cooperative stage of the biform game $\left({{\text{N}},{v_{f}},{\text{X}}}\right)$. $\left({N,\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$ is called the non-cooperative game derived by $\left({N,{v_{f}},X}\right)$.
@@ 2.3 Distribution of Cooperative Outcomes and Biform Game Solutions
Definition 2.3 If the strategy combination $x\in X$ is a Nash equilibrium of the non-cooperative game $\left({N,\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$ derived from the biform game, that is, $x$ satisfies
@@
Then $x$ is called the solution of the biform game $\left({N,{v_{f}},X}\right)$.
Remark 2.3 From the definition of the solution of the biform game, the allocation function $a_{i}^{{v_{f}}}$ generated in the cooperation stage plays a crucial role.
@@
We call $a^{{v_{f}}}$} the Shapley allocation function of $v_{f}$}, abbreviated as ${a^{{v_{f}}}}=Sh\left({{v_{f}}}\right)$. It is not difficult to verify that if the marginal function $m_{i}^{{v_{f}}}$ is payment-dominant, then the allocation function $Sh\left({{v_{f}}}\right)$ is also marginalist. In fact, if ${f_{i}}\left(x\right)\leqslant{f_{i}}\left(y\right)$ holds for every $i\in N$ and any $x,y\in X$, it is equivalent to $m_{i}^{{v_{f}}}\left({S,x}\right)\leqslant m_{i}^{{v_{f}}}\left({S,y}\right)$ holding for every $S\subset N$, $i\in N$ and any $x,y\in X$, which is equivalent to $S{h_{i}}\left({v\left({\cdot,x}\right)}\right)\leqslant S{h_{i}}\left({v\left({\cdot,y}\right)}\right)\;$holding for every $i\in N$ and any $x,y\in X$ holds, that is, ${\text{if\;and\;only\;if\;}}a_{i}^{{v_{f}}}\left(x\right)=S{h_{i}}\left({v\left({\cdot,x}\right)}\right)\leqslant S{h_{i}}\left({v\left({\cdot,y}\right)}\right)=a_{i}^{{v_{f}}}\left(y\right)$ holding for every $i\in N$ and any $x,y\in X$. Hence, $Sh\left({{v_{f}}}\right)={a^{{v_{f}}}}$ is marginalist.
(2) If there is no additional benefit, that is, the remainder $\delta=0$, ${v_{f}}={u_{f}}$. In this case, $a^{{u_{f}}}$} is the Shapley allocation function of $u_{f}$}, that is
@@
Below, we will analyze the profound impact of the two distribution methods of marginalism and egalitarianism on the solution of biform games.
Proposition 2.1 If the allocation function $a^{{v_{f}}}$} is marginalist, then a strategy combination $x$ is a solution of $\left({N,{v_{f}},X}\right)$ if and only if $x$ is a Nash equilibrium of $\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$.

### v2

# The Allocation of Interests and The Maximization of Each Player’s Benenfits:Transforming Strategic Games into Biform Games and Applications

@@ Abstract
As Aumann stated, cooperation and non-cooperation are different ways of viewing the same game, with the main difference being whether players can reach a binding cooperative agreement. In the real world, many games often coexist competition and cooperation. Based on the above reasons, we propose a method to transform strategic games into a biform game model, which retains the characteristics of cooperative games while considering the ultimate goal of players to maximize their own interests. Among them,the allocation function play a crucial role as a protocol on convention. Furthermore, based on this game model, we analyze the impact of two different distribution methods, namely marginalism and egalitarianism, on the game results. As an application, we analyze how food producers seek maximum profits through cooperative pricing.
@@ 1 Introduction
In this paper, we will propose a new research framework for any strategic game. It is a game model that takes into account both the cooperative and non-cooperative behaviors of the players. This framework overcomes the defects encountered in directly present the characteristic functions.To further verify the practical value of the proposed framework, As applications, the choice of green technology in the market competition model will be discussed. It is evident from these examples that the new research framework can effectively break through the dilemmas embedded in them.
This paper mainly includes two parts: the first part is to establish a research method of transforming from a strategic game (players can choose a cooperative game or a non-cooperative game) to a biform game; in the second part, we illustrate the method and sentencing game model proposed in the first part through two examples, and utilize this conversion method and biform model to analyze how food producers seek maximum profits through cooperative pricing.
@@
Let’s turn our attention to cooperative game theory, and specifically focous on the distribution of cooperation outcomes.The most important thing in cooperative game theory is the solution to the cooperative game, which is to answer the question: When we assume that the result of cooperation is the benefit of the grand coalition, how should the result of cooperation be distributed among the players in a reasonable way? The rationality mentioned here includes fairness, the enthusiasm to encourage players to form alliances, and a certain degree of stability (the coalition is not easily destroyed). From a large body of literature, it is not difficult to find that the solutions of cooperative games are inseparable from the two logics of marginalism and egalitarianism. The Shapley value is a typical example of marginalism, while the equal division rule is a typical example of egalitarianism [14, 15, 16, 17, 18]. The Shapley value does not allocate anything to non-productive players (those who do not make a contribution), while the equal allocation rule does not consider the contribution rate of players to the coalition’s revenue. However, neither of these two methods is completely consistent with the form of resource or outcome distribution in practical problems. When distributing the benefits of cooperation, we not only give incentives to players with high contributions, but also pay attention to the common sharing of the benefits of cooperation. What impact will marginalism and egalitarianism have on the final result in the distribution? How should our distribution method seek a balance between the two for specific problems? Or, what kind of distribution plan should we formulate in order to achieve the goal? These are the questions that the first part of this paper attempts to answer. This part will systematically study the impact of the distribution of coalition benefits based on the biform game derived from the strategic game, and explore a method to analyze how the distribution of coalition utility affects strategic behavior.
@@ 2.1 Problem Statement
Suppose the strategic game is $\left({N,X,f}\right)$. Here $N=\left\{{1,\cdots,n}\right\}$ is the set of players, $X_{i}$} represents the strategy set of participant $i$, $X=\mathop{\prod}\nolimits_{i\in N}{X_{i}}$ is the strategy profile set, ${x_{i}}\in{X_{i}}$ represents the strategy selected by participant $i$, $\left({{x_{i}},{x_{-i}}}\right)\in{\text{X}}$ represents the strategy profile formed after each participant $i$ selects $x_{i}$}, and $x_{-i}$} represents the strategy selection component of all players except $i$. $f=\left({{f_{1}},\cdots,{f_{n}}}\right)$ is a vector of payoff functions, where ${f_{i}}:X\to R$ is the payoff function of player $i$. In particular, when $N=\left\{{1,2}\right\}$ and ${X_{1}}={S_{1}}$ and ${X_{2}}={S_{2}}$ are finite sets, then the two-person finite game $\left({\left\{{1,2}\right\},\left({{S_{i}}}\right),\left({{f_{i}}}\right)}\right)$ can be represented by a table, that is, the two payoff matrices are combined into one table.
When a strategic game is considered a non-cooperative game, we still use its notation, but assign the strategy set and payoff function to each participant, as, $\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$. In this scenario, each participant seeks to maximize their own payoff. A state of equilibrium, known as the Nash equilibrium the central solution concept in non-cooperative game theory is reached when no player can obtain a higher payoff by unilaterally changing their strategy.
@@
In non-zero-sum situations, we have ${f_{1}}\left({x_{1}^{*},x_{2}^{*}}\right)>v\left({\left\{1\right\}}\right)$ and ${f_{2}}\left({x_{1}^{*},x_{2}^{*}}\right)>v\left({\left\{2\right\}}\right)$. For example, the Battle of the Sexes game, the Hawk-Dove game, and the Wise Pigs game (also known as the free-rider game) all satisfy this property. Note that ${f_{1}}\left({x_{1}^{*},x_{2}^{*}}\right)$ and ${f_{2}}\left({x_{1}^{*},x_{2}^{*}}\right)$ are the payoffs under non-cooperation, while $v\left({\left\{1\right\}}\right)$ and $v\left({\left\{2\right\}}\right)$ are the payoffs under taking into account the possibility of cooperation. Considering the opponent’s confrontation actually reduces the payoffs of the players, which seems logically unreasonable. Therefore, it is unreasonable to assume that players other than $S$ choose a confrontational strategy.
@@
Furthermore, transforming strategic games into cooperative games focuses on cooperation between players, but fails to fully reflect each player’s pursuit of self-interest. To overcome these limitations, we propose a biform game model derived from strategic games. This model assigns a coalition value to each strategy profile, regardless of the strategies chosen by players within or outside the coalition. Cooperation within the coalition manifests itself primarily through utility transfer and the potential generation of external benefits.
@@ 2.2 Model Building
Similar to expressions (1), (2) and (3), ${u_{f}}\left({S,x}\right)$ represents the payoff of coalition $S$ at the strategy profile $x$, referred to as the value of coalition $S$ at $x$, which is the total payoff of the players in coalition $S$.
As mentioned earlier, the characteristic function defined by (1), (2) and (3) must be either the maximum or minimum value on $X_{S}$} or the corresponding $X_{-S}$}. By defining the coalition value according to (4), there is no need to worry about whether to maximize or minimize it, as with $v$, $v_{H}$}, and $v_{M}$}. A way to circumvent this problem is to give, for each strategy profile $x\in X$, the value ${u_{f}}\left({S,x}\right)$ of the coalition $S$ at $x$, then, ${u_{f}}\left({\cdot,x}\right):{2^{X}}\to R$ is the characteristic function of the strategy profile $x$, or the coalition function associated with the strategy profile $x$. For each $x\in X$, ($N,{u_{f}}\left({\cdot,x}\right)$) constitutes a cooperative game.
According to the biform game model proposed by Brandenburger and Stuart [4], cooperative games with strategy profiles $x$ constitute biform games. Biform games first appeared in concrete examples ([2] and Hart and Moore [3]), and the concept of biform games and their solutions was first introduced by Brandenburger and Stuart [4].
@@
Here $\delta\left({S,x}\right)\geqslant 0$ refers to the additional benefits generated by cooperation, beyond the cumulative gains of the players. in some cases, coalitions can bring additional benefits, such as collaborative effects, government subsidies, and cost savings after supply chain members form a coalition, as well as logistics and warehousing costs. Therefore,
@@
We still call ${v_{f}}:{2^{N}}\times X\mapsto R$ the characteristic function with the strategy profile $x$, or the coalition function with the strategy profile $x$. In this case, $u_{f}$} becomes a special case of $v_{f}$}, that is, the case where the additional benefit $\delta\left({\cdot,x}\right)=0$.
Furthermore, considering that when players cooperate, in addition to the utility transfer and collaborative effects between players, another possibility is to restrict the scope of strategy choices through agreement and understandings. From the perspective of the strategy profile set $X$, this means adding constraints or limiting strategy choices to a subset $X^{C}$} of $X$. $X^{C}$} is called the collaboration set of the strategic game $\left({N,X,f}\right)$, or simply the collaboration set. Based on the above analysis, the following definition is given.
Definition 2.1 For a given strategic game $\left({N,X,f}\right)$, ${v_{f}}:{2^{N}}\times X\mapsto R$ is the characteristic function with the strategy profile $x$, ${X^{C}}\subset X$ is the cooperation set, then $\left({N,{v_{f}},{X^{C}}}\right)$ constitutes a biform game, which is called a biform game derived from $\left({N,X,f}\right)$.
Remark 2.1 (1) For every $x\in X$, $\left({N,{v_{f}}\left({\cdot,x}\right)}\right)$ is the cooperative game with the strategy profile $x$, also called the cooperative game at $x$ derived from $\left({N,X,f}\right)$.
@@
(3) ${X^{C}}\subset X$ represents a mutual agreement among players to appropriately limit the range of strategic options, including government intervention in prices and production capacity under certain circumstances. A major role of the Organization of the Petroleum Exporting Countries (OPEC) is to restrict oil prices or production for its members. Recently, the state’s efforts to combat involution in certain industries have also involved interventions in production capacity and prices.
For the biform game $\left({N,{v_{f}},X}\right)$ derived from the strategic game $\left({N,X,f}\right)$, the first stage is a non-cooperative game, primarily describing the strategic choices of the players. This involves the interactive process of each participant selecting different strategies $x_{i}$} (forming a strategy profile $x$) to maximize their payoff. This also reflects the mutual influence of strategy choices as players pursue maximum payoff. The second stage is concerned with each cooperative game $\left({N,{v_{f}}\left({\cdot,x}\right)}\right)$, which reflects the formation of alliances among players and the distribution of cooperative gains among them. It determines the payoff value each participant receives when a strategy profile is realized, and also serves as the payoff function for the first-stage non-cooperative game. Transforming a strategic game into a biform game represents a novel analytical approach. It not only addresses the shortcomings of directly converting a strategic game into a cooperative game such as relying on max-min methods but also constitutes a new attempt to consider competition and cooperation simultaneously.
@@
Let $N=\left\{{1,2}\right\}$ denote the set of players (two herders), ${X_{1}}={X_{2}}=\left\{{C,NC}\right\}$ denote their strategy set, $X={X_{1}}\times{X_{2}}$ denote the strategy profile set, ${f_{1}}\left(x\right),{f_{2}}\left(x\right)$ denote the payoff functions of herders 1 and 2 respectively, and each $f\left(x\right)=\left({{f_{1}}\left(x\right),{f_{2}}\left(x\right)}\right)$ denote a payoff vector, where
@@
2.3 Distribution of Cooperative Outcomes and Biform Game Solutions
[…]
@@ 2.3 Distribution of Cooperative Outcomes and Biform Game Solutions
For each $i\in N$ and $S\subset N$, if $i\in S$, set $m_{i}^{{v_{f}}}=0$. Then $m_{i}^{{v_{f}}}:{2^{N}}\times X\to R$ is called the marginal function of player $i\in S$ with respect to $v_{f}$}. Let $m^{{v_{f}}}(S,X)=(m_{1}^{{v_{f}}},m_{2}^{{v_{f}}},...,m_{n}^{{v_{f}}})$ . $m^{{v_{f}}}:{2^{N}}\times X\to R^{N}$ is called the marginal function respect to $v_{f}$}.
Definition 2.2 Let $\left({N,{v_{f}},X}\right)$ be a biform game derived from $\left({N,X,f}\right)$. The marginal function $m^{{v_{f}}}$ is called payoff-dominant if for every $i\in N$ and $x,y\in X$, ${f_{i}}\left(x\right)>{f_{i}}\left(y\right)$, it follows that $m_{i}^{{u_{f}}}\left({S,x}\right)>m_{i}^{{u_{f}}}\left({S,y}\right)$ for all $S\subset N$.
Remark 2.2 If for every $i\in N$ and $x,y\in X$, ${f_{i}}\left(x\right)>{f_{i}}\left(y\right)$, it follows that $m_{i}^{\delta}\left({S,x}\right)>m_{i}^{\delta}\left({S,y}\right)$, then $m_{i}^{{v_{f}}}$ is payment-dominated. This means that the larger the payoff to the participant, the larger the value of the additional benefit distributed. In particular, if $\delta\left({\cdot,x}\right)=c$, it is clear that $m_{i}^{{v_{f}}}$ is payment-dominated, and thus it is clear that $m^{{u_{f}}}$ is payment-dominated.
Before presenting the solution to the biform game, Let us start addressing a critical issue: the distribution of coalition payoffs during the cooperative phase. This determines the important properties of the biform game solution. When returning to the first stage of the non-cooperative game, the Nash equilibrium will be primarily influenced by the strategic choices of the players, whose goal is to maximize their respective payoffs, which are determined by the allocation function. In numerous practical problems, the allocation function is typically determined by either a formalized protocol-with clear definitions of the rights and obligations of all parties-or well-established conventions validated by long-term industry practice. This determination mechanism ensures the rationality and enforceability of distribution within a coalition.
Suppose the vector ${a^{{v_{f}}}}\left(x\right)=\left({a_{1}^{{v_{f}}}\left(x\right),\cdots,a_{n}^{{v_{f}}}\left(x\right)}\right)$ is the solution of the cooperative stage $\left({N,{v_{f}}\left({\cdot,x}\right)}\right)$ of the cooperative game. Each $a_{i}^{{v_{f}}}\left(x\right)$ is the distribution of the grand coalition payoff ${v_{f}}\left({N,x}\right)$ to the payoff of participant $i$. We call the vector-valued function ${a^{{v_{f}}}}:X\mapsto{R^{n}}$ an allocation function of $v_{f}$}. The function $a_{i}^{{v_{f}}}\left(x\right)$ gives a new payoff to player $i$, which replaces the initial payoff value ${f_{i}}\left(x\right)$ in the strategic game. The non-cooperative game $\left({N,\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$ constitutes the non-cooperative stage of the biform game $\left({N,{v_{f}},X}\right)$. $\left({N,\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$ is called the non-cooperative game derived from $\left({N,{v_{f}},X}\right)$.
@@
Definition 2.3 If the strategy profile $x\in X$ is a Nash equilibrium of the non-cooperative game $\left({N,\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$ derived from the biform game, that is, $x$ satisfies
@@
Then $x$ is called the solution of the biform game $\left({N,{v_{f}},X}\right)$, also called the soution of the biform game $\left({N,X,f}\right)$.
Remark 2.3 According to the definition of the solution in the biform game, the allocation function $a_{i}^{{v_{f}}}$ generated in the cooperation stage is crucial. That is, the solution is determined by this allocation function.
@@
We call $a^{{v_{f}}}$} the Shapley allocation function of $v_{f}$}, abbreviated as ${a^{{v_{f}}}}=Sh\left({{v_{f}}}\right)$. It is not difficult to verify that if the marginal function $m^{{v_{f}}}$ is payment-dominant, then the allocation function $Sh\left({{v_{f}}}\right)$ is also marginalist. In fact, if ${f_{i}}\left(x\right)\leqslant{f_{i}}\left(y\right)$ holds for every $i\in N$ and any $x,y\in X$, it is equivalent to $m_{i}^{{v_{f}}}\left({S,x}\right)\leqslant m_{i}^{{v_{f}}}\left({S,y}\right)$ holding for every $S\subset N$, $i\in N$ and any $x,y\in X$, which is equivalent to $S{h_{i}}\left({v\left({\cdot,x}\right)}\right)\leqslant S{h_{i}}\left({v\left({\cdot,y}\right)}\right)\;$holding for every $i\in N$ and any $x,y\in X$ holds, that is, ${\text{if\;and\;only\;if\;}}a_{i}^{{v_{f}}}\left(x\right)=S{h_{i}}\left({v\left({\cdot,x}\right)}\right)\leqslant S{h_{i}}\left({v\left({\cdot,y}\right)}\right)=a_{i}^{{v_{f}}}\left(y\right)$ holding for every $i\in N$ and any $x,y\in X$. Hence, $Sh\left({{v_{f}}}\right)={a^{{v_{f}}}}$ is marginalist.
(2)Distribution being marginalist means that within $v_{f}(S,x)=\mathop{\sum}\nolimits_{i\in S}f_{i}(x)+\delta(x)$ those whose contributions are greater shall be allocated a larger portion of the distribution. If there is no additional benefit, that is, the remainder $\delta=0$, ${v_{f}}={u_{f}}$. In this case, $a^{{u_{f}}}$} is the Shapley allocation function of $u_{f}$}, that is
@@
The allocation function is of vital importance in solving the biform game $(N,v_{f},X)$, and how to determine it depends on the actual problems. We will analyze the profound impact of the two distribution methods of marginalism and egalitarianism on the solution of biform games.
Proposition 2.1 If the allocation function $a^{{v_{f}}}$} is marginalist, then a strategy profile $x$ is a solution of $\left({N,{v_{f}},X}\right)$ if and only if $x$ is a Nash equilibrium of $\left({N,\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$.